# Adjusted Connections I: Differential Cocycles for Principal Groupoid Bundles with Connection

### v1

@@ 5.1 Atiyah algebroids from inner action groupoids
We will generalize this picture in two ways. First of all we note that for the discussion of groupoid bundles with connections404040as well as for a discussion of higher principal bundles with connections, cf. [FJKSa] a splitting as vector bundles unfortunately does not reproduce the gauge transformations. Instead, one has to regard the vector bundles in the sequence as Lie algebroids (e.g. $\mathrm{T}X$ is here the tangent algebroid) and split in this category. This, in turn, only leads to flat connections, and we will have to extend this picture. Just as in the local description of Section 3, we will have to construct general connections as restricted higher flat connections. This construction will make use of adjustment Lie groupoids.
Second, for performing concrete computations, e.g. in a field theory having groupoid connections as kinematical data, it is much more convenient to switch to a local perspective in terms of cocycles and local connection forms. In this perspective, we replace the manifold $X$ by the Morita-equivalent Čech groupoid $\check{\mathscr{C}}(\sigma)$ of a surjective submersion $\sigma:Y\rightarrow X$.
@@
for $\nu\in\mathfrak{g}[1]$, $g\in\mathsf{G}$ and $\gamma\in\mathrm{T}[1]_{g}\mathsf{G}$, where $\mathbbm{1}$ is the unit element of $\mathsf{G}$, and differential given by the Chevalley–Eilenberg differential on $\mathfrak{g}[1]$ and the de Rham differential on $\mathsf{G}$.
@@
Consider a general cone $\mathcal{C}=(\mathcal{C}_{1}\rightrightarrows\mathcal{C}_{0})$ and fit it into the diagram
@@ 5.2 Principal bundles with connections as dg-Lie groupoid morphisms
As done in the local description Section 3, we now obtain general connections on principal bundles as higher connections with restricted gauge transformations. The relevant classifying space here is precisely the adjustment groupoid. We start by specializing the adjustment groupoid constructed in Section 4.3 to the case $\mathscr{G}=\mathsf{B}\mathsf{G}$. Note that the connection on $\mathsf{B}\mathsf{G}$ is necessarily trivial.
@@
$\begin{gathered}\mathsf{t}(\nu,\bar{\nu};g,\gamma)=(\nu,\bar{\nu})~{},~{}~{}~{}\mathsf{s}(\nu,\bar{\nu};g,\gamma)=\left(g^{-1}\nu g+g^{-1}\gamma,g^{-1}\bar{\nu}g\right)~{},\\ (\nu,\bar{\nu};g_{1},\gamma_{1})\circ\left(g_{1}^{-1}\nu g_{1}+g_{1}^{-1}\gamma_{1},g^{-1}_{1}\bar{\nu}g_{1};g_{2},\gamma_{2}\right)=(\nu,\bar{\nu};g_{1}g_{2},g_{1}\gamma_{2}+\gamma_{1}g_{2})\\ \mathsf{e}(\nu,\bar{\nu})=(\nu,\bar{\nu};\mathbbm{1},0)~{},~{}~{}~{}(\nu,\bar{\nu};g,\gamma)^{-1}=\left(g^{-1}\nu g+g^{-1}\gamma,g^{-1}\bar{\nu}g;g^{-1},-g^{-1}\gamma g^{-1}\right)\end{gathered}$ (5.19)
@@
[…]
For uniqueness, we use the fact that $\mathsf{t}$ is a submersion of graded manifolds, which allows us to write the morphism $\mathscr{A}_{1}$ of a general adjustment groupoid of $\mathsf{B}\mathsf{G}$ as $\mathscr{A}_{1}=\mathrm{T}[1]\mathfrak{g}[1]\times\mathcal{R}$ for $\mathcal{R}$ some dg-manifold. The pullback condition of pre-adjustment structures
$\mathfrak{g}[1]\times\mathrm{T}[1]\mathsf{G}=\mathfrak{g}[1]\mathbin{{}_{(\mathrm{id},0)}\times_{\mathsf{t}}}(\mathrm{T}[1]\mathfrak{g}[1]\times\mathcal{R})$ (5.21)
implies that $\mathcal{R}=\mathrm{T}[1]\mathsf{G}$. The freedom here is precisely the isomorphism of $\mathscr{A}_{1}$ preserving the embedding of $\mathfrak{g}[1]$. ∎
@@ 6 Principal groupoid bundles with connections
We have now everything at hand to generalize the discussion of the previous section to arbitrary principal groupoid bundles.
@@ 6.2 Flat and adjusted connections on principal groupoid bundles
If the adjustment groupoid is of the form $\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}$ of Theorem 4.14, then infinitesimal gauge transformations act as
@@
The fact that the gauge transformations are of this form follows directly from the proof of Theorem 4.14, together with the derivation of adjusted gauge transformations in Section 3.3. ∎

### v3

@@ 5.1 Atiyah algebroids from inner action groupoids
We will generalize this picture in two ways. First of all we note that for the discussion of groupoid bundles with connections424242as well as for a discussion of higher principal bundles with connections, cf. [FJKSa] a splitting as vector bundles unfortunately does not reproduce the gauge transformations. Instead, one has to regard the vector bundles in the sequence as Lie algebroids (e.g. $\mathrm{T}X$ is here the tangent algebroid) and split in this category. This, in turn, only leads to flat connections, and we will have to extend this picture. Just as in the local description of Section 3, we will have to construct general connections as restricted higher flat connections. This construction will make use of adjustment Lie groupoids.
Second, for performing concrete computations, e.g. in a (physical) field theory having groupoid connections as kinematical data, it is much more convenient to switch to a local perspective in terms of cocycles and local connection forms. In this picture, we replace the manifold $X$ by the Morita-equivalent Čech groupoid $\check{\mathscr{C}}(\sigma)$ of a surjective submersion $\sigma:Y\rightarrow X$.
@@
for $\nu\in\mathfrak{g}[1]$, $g\in\mathsf{G}$ and $\gamma\in\mathrm{T}[1]_{g}\mathsf{G}$, where $\mathbbm{1}$ is the unit element of $\mathsf{G}$, and the differential is given by the Chevalley–Eilenberg differential on $\mathfrak{g}[1]$ and the de Rham differential on $\mathsf{G}$.
@@
Consider a general cone $\mathcal{C}=(\mathcal{C}_{1}\rightrightarrows\mathcal{C}_{0})$, and fit it into the diagram
@@
Remark 5.3.
In the diagram (5.9), the weak pullback coincide with the strict pullback which can be proved abstractly. However, we omit this not to complicate the discussion.
@@ 5.2 Principal bundles with connections as dg-Lie groupoid morphisms
As done in the local description Section 3, we now obtain general connections on principal bundles as higher connections with restricted gauge transformations. The relevant classifying space here is precisely the adjustment groupoid. We start by specializing the adjustment groupoid of Section 4.3 to the case $\mathscr{G}=\mathsf{B}\mathsf{G}$. Note that the connection on $\mathsf{B}\mathsf{G}$ is necessarily trivial.
@@
$\begin{gathered}\mathsf{t}(\nu,\bar{\nu};g,\gamma)=(\nu,\bar{\nu})~{},~{}~{}~{}\mathsf{s}(\nu,\bar{\nu};g,\gamma)=\left(g^{-1}\nu g+g^{-1}\gamma,g^{-1}\bar{\nu}g\right)~{},\\ (\nu,\bar{\nu};g_{1},\gamma_{1})\circ\left(g_{1}^{-1}\nu g_{1}+g_{1}^{-1}\gamma_{1},g^{-1}_{1}\bar{\nu}g_{1};g_{2},\gamma_{2}\right)=(\nu,\bar{\nu};g_{1}g_{2},g_{1}\gamma_{2}+\gamma_{1}g_{2})~{},\\ \mathsf{e}(\nu,\bar{\nu})=(\nu,\bar{\nu};\mathbbm{1},0)~{},~{}~{}~{}(\nu,\bar{\nu};g,\gamma)^{-1}=\left(g^{-1}\nu g+g^{-1}\gamma,g^{-1}\bar{\nu}g;g^{-1},-g^{-1}\gamma g^{-1}\right)\end{gathered}$ (5.19)
@@
This is the specialization of Theorem 4.8. Once it is noted that there is a unique Cartan connection (which is trivial) on a Lie group, the proof is complete.
@@ 6 Principal groupoid bundles with connections
We now have everything at hand to generalize the discussion of the previous section to arbitrary principal groupoid bundles.
@@ 6.2 Flat and adjusted connections on principal groupoid bundles
Infinitesimal gauge transformations act as
@@
The fact that the gauge transformations are of this form follows directly from Lemma 6.6 and Remark 3.15, together with the derivation of adjusted gauge transformations in Section 3.3. ∎
Lemma 6.6 (Adjustment groupoid).
For any Cartan Connection $\mathcal{H}$ on a Lie groupoid $\mathscr{G}$, there is an isomorphism of presheaves consist of dg-maps
$\mathcal{N}\mapsto\underline{\mathsf{hom}}^{\rm red}_{\nabla_{\mathcal{H}}}(\mathcal{N},\mathrm{T}[1]\mathfrak{g}[1])~{}~{}~{}\mbox{and}~{}~{}~{}\mathcal{N}\mapsto\mathsf{Lie}\left(\mathsf{dgGrpd}\left(\mathcal{N}\rightrightarrows\mathcal{N},\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}\right)\right)~{},$ (6.4)
where $\nabla_{\mathcal{H}}$ is the Cartan connection on the Lie algebroid $\mathfrak{g}$ induced by $\mathcal{H}$.
Proof.
This requires a short characterization of the first presheaf and a rather lengthy computation for the second presheaf.
[…]
  • a topological map
    $|f_{\flat}|:|\mathcal{X}|\times|\mathcal{N}|\rightarrow M$ (6.5)
  • and two maps of modules over $C^{\infty}(M)$,
    $f^{\sharp}_{\flat 0}:$ $\Gamma(\mathfrak{g}[1]^{*})$ $\rightarrow|f|_{*}\left(C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})_{1}\right)~{},$ (6.6)
    $f^{\sharp}_{\flat 1}:$ $\Gamma(\mathfrak{g}[1]^{*})$ $\rightarrow|f|_{*}\left(C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}\right)~{},$
    subject to conditions explained below and listed in (6.8).
The additional conditions ensure that $f^{\sharp}\circ\mathsf{p}_{\rm curv}^{\sharp}$ only takes values in $C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})$. The lift $f$ can be explicitly calculated by imposing that $f$ is a dg-map. In the local coordinates $(m^{a},\bar{m}^{a},\xi^{\alpha},\bar{\xi}^{\alpha})$ and the structure functions introduced in (2.55), $f$ is given by $|f|=|f_{\flat}|$ and
$\begin{split}f^{\sharp}(m^{a})&=f^{\sharp}_{\flat}(m^{a})\\ &=|f_{\flat}|^{*}(m^{a})~{},\\ f^{\sharp}(\xi^{\alpha})&=f^{\sharp}_{\flat}(\xi^{\alpha})\\ &=f^{\sharp}_{\flat 0}(\xi^{\alpha})+f^{\sharp}_{\flat 1}(\xi^{\alpha})~{},\\ f^{\sharp}(\bar{m}^{a})&=(\mathrm{d}_{\mathcal{X}}+\mathrm{d}_{\mathcal{N}})f^{\sharp}_{\flat}(m^{a})-f^{\sharp}_{\flat}(\mathrm{d}_{\mathfrak{g}[1]}(m^{a}))\\ &=(\mathrm{d}_{\mathcal{X}}+\mathrm{d}_{\mathcal{N}})f^{\sharp}_{\flat}(m^{a})-|f_{\flat}|^{*}\mathtt{r}^{a}_{\alpha}\left(\tilde{f}^{\sharp}_{\flat 0}(\xi^{\alpha})+f^{\sharp}_{\flat 1}(\xi^{\alpha})\right)~{},\\ f^{\sharp}(\bar{\xi}^{\alpha})&=(\mathrm{d}_{\mathcal{X}}+\mathrm{d}_{\mathcal{N}})f^{\sharp}_{\flat}(\xi^{\alpha})-f^{\sharp}_{\flat}(\mathrm{d}_{\mathfrak{g}[1]}(\xi^{\alpha}))~{}.\end{split}$ (6.7)
For $f^{\sharp}\circ\mathsf{p}_{\rm curv}^{\sharp}$ to only take values in $C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})$, we need to impose the following relations:
$\begin{gathered}\mathsf{d}_{\mathcal{X}}|f|^{*}h=\left(f^{\sharp}_{1}\circ\rho^{*}\right)(\mathrm{d}h)~{},\\ \mathsf{d}_{\mathcal{X}}f_{0}^{\sharp}\left(\xi^{\alpha}\right)+|f|^{*}\left(\mathtt{f}_{\beta\gamma}^{\alpha}\right)f^{\sharp}_{1}\left(\xi^{\beta}\right)f^{\sharp}_{0}\left(\xi^{\gamma}\right)+|f|^{*}\left(\omega_{a\beta}^{\alpha}\right)f^{\sharp}\left(\bar{m}^{a}\right)\tilde{f}^{\sharp}_{1}\left(\xi^{\beta}\right)+\mathsf{d}_{\mathcal{N}}f^{\sharp}_{1}\left(\xi^{\alpha}\right)=0~{},\\ \mathsf{d}_{\mathcal{X}}f^{\sharp}_{1}\left(\xi^{\alpha}\right)=-\tfrac{1}{2}|f|^{*}\left(\mathsf{f}_{\beta\gamma}^{\alpha}\right)f^{\sharp}_{1}\left(\xi^{\beta}\right)f^{\sharp}_{1}\left(\xi^{\gamma}\right)\end{gathered}$ (6.8)
for all $h\in C^{\infty}(M)$.
that factor through another groupoid:
where the topological part $|\mathsf{p}_{0}|$ of $\mathsf{p}_{0}$ is identity, $\mathsf{p}_{0}^{\sharp}$ is the evident inclusion of degree 00 elements in the local ring of functions, and $\check{F}$ is some morphism of graded manifolds. In the following, we will show that such functors are given by
  • a topological map
    $|F|:|\mathcal{X}|\times|\mathcal{N}|\to M$ (6.12)
  • and two maps of modules over $C^{\infty}(M)$
    $(\bar{F}_{0})^{\sharp}:$ $\Gamma(\mathfrak{g}[1])^{*}$ $\to|F|_{*}\left(C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})_{1}\right)~{},$ (6.13)
    $(\bar{F}_{1})^{\sharp}:$ $\Gamma(\mathfrak{g}[1])^{*}$ $\to|F|_{*}\left(C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}\right)~{},$
A general functor $F$ as in (6.10) is given by a pair of dg-maps between morphisms and objects,
$F_{1}:$ $\mathcal{X}\times\Theta\times\Theta\times\mathcal{N}$ $\to\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}~{},$ (6.14)
$F_{0}:$ $\mathcal{X}\times\Theta\times\mathcal{N}$ $\to\mathrm{T}[1]\mathfrak{g}[1]~{},$
satisfying functoriality conditions. The dg-map $F_{0}:\mathcal{X}\times\Theta\times\mathcal{N}\rightarrow\mathrm{T}[1]\mathfrak{g}[1]$ is determined by uniquely lifting a graded map $F_{\flat 0}:\mathcal{X}\times\Theta\times\mathcal{N}\rightarrow\mathfrak{g}[1]$ which is given by a topological map $|F_{\flat 0}|=|F|$ as well as maps
$\left(F_{\flat 0}^{\sharp}\right)_{0}:$ $C^{\infty}(M)$ $\to|F|_{*}\left(C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})_{0}~{}\oplus~{}\theta C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}\right)~{},$ (6.15)
$\left(F_{\flat 0}^{\sharp}\right)_{1}:$ $\Gamma(\mathfrak{g}[1])^{*}$ $\to|F|_{*}\left(C^{\infty}(\mathcal{X})_{0}\otimes C^{\infty}(\mathcal{N})_{1}~{}\oplus~{}\theta C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{1}\right)~{},$
where $(F_{\flat 0}^{\sharp})_{0}$ is a ring homeomorphism, $(F_{\flat 0}^{\sharp})_{1}$ is a morphism of modules over $C^{\infty}(M)$ and $\theta$ is the generator of $C^{\infty}(\Theta)$ with $\mathrm{d}_{\Theta}\theta=1$. The map $(F_{\flat 0}^{\sharp})_{0}$ is calculated similar to (6.7). Because the images of $F_{\flat 0}^{\sharp}$ on the “curvature generators” $\bar{m}^{a}$ and $\bar{\xi}^{\alpha}$ do not belong to
for all $h\in C^{\infty}(M)$ and $s\in\Gamma(\mathfrak{g}[1]^{*})$.
Compatibility with the source maps then reads as
By our usual argument from Remark A.1, the map $F_{1}^{\prime}$ is given by a lift of a graded map $F_{1\flat}^{\prime}:\mathcal{X}\times\Theta\times\Theta\times\mathcal{N}\to\mathsf{G}$ and (6.18) translates to $\mathsf{t}\circ F_{1\flat}^{\prime}=\mathsf{p}_{M}\circ\psi\circ F_{0}\circ\mathsf{t}$, where $\mathsf{p}_{M}$ is again the projection to the underlying manifold $M$ of $\mathrm{T}[1]M$. Furthermore, as $|\mathscr{P}\mathsf{air}(\Theta)|$ is a trivial Lie groupoid and $|\mathscr{A}^{\mathsf{W}}_{\mathcal{H}}|=(\mathsf{G}\rightrightarrows M)$, the condition $\mathsf{e}\circ F_{0}=F_{1}\circ\mathsf{e}$ implies
$|F_{1}^{\prime}|=|F_{1}|=\mathsf{e}\circ|F_{0}|=\mathsf{e}\circ|F|~{}.$ (6.20)
[…]
and satisfy the identities
[…]
The map $F_{1\flat}^{\prime\prime}$ is now defined by $|F^{\prime\prime}_{1\flat}|=|F|$ and a derivation
$H\colon C^{\infty}(G)\to\left(\mathsf{e}\circ|F|\right)_{*}\left(C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}\right)$ (6.25)
according to
$\begin{gathered}(F^{\prime\prime}_{1\flat})^{\sharp}(g)=\left(\mathsf{e}\circ|F|\right)^{*}g-\theta H(g)~{}\in~{}C^{\infty}(|\mathcal{X}|\times|\mathcal{N}|)~{}\oplus~{}\theta C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}\end{gathered}$ (6.26)
for $g\in C^{\infty}(\mathsf{G})$. As argued in [Li14, Prop. 8.21], any such derivation is given by the de Rham differential followed by a morphism of modules
$H^{\prime}\colon\Omega^{1}(\mathsf{G})\to\left(\mathsf{e}\circ|F|\right)_{*}\left(C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}\right)~{},$ (6.27)
$H^{\prime}\left(\mathsf{t}^{*}\varpi\right)=0$ (6.28)
$H^{\prime\prime}\colon\Gamma(\ker(\mathrm{d}\mathsf{t}))^{*}\to\left(\mathsf{e}\circ|F|\right)_{*}\left(C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}\right)$ (6.29)
and $\mathsf{i}^{*}$ is induced by $\mathsf{i}\colon\mathsf{Ker}(\mathrm{d}\mathsf{t})\to\mathrm{T}\mathsf{G}$. Finally, $H^{\prime\prime}$ is a map of sheaves, and therefore it factors uniquely through pullbacks if the topological part does. $H^{\prime\prime}$ is uniquely determined by
$H^{\prime\prime\prime}\colon\mathsf{e}^{*}\Gamma(\ker(\mathrm{d}\mathsf{t}))^{*}=\Gamma(\mathfrak{g}[1])^{*}\to|F|_{\star}\left(C^{\infty}(\mathcal{X})_{1}\otimes C^{\infty}(\mathcal{N})_{0}\right)~{}.$ (6.30)
Therefore, we can identify $H^{\prime\prime\prime}$ with $(\bar{F}_{1})^{\sharp}$.
$\begin{gathered}|F^{\prime}_{1\flat}|:|\mathcal{X}|\times|\mathcal{N}|\to\mathsf{G}~{}~{}~{}\mbox{with}~{}~{}~{}|F^{\prime}_{1\flat}|=\mathsf{e}\circ|F|~{},\\ (F^{\prime}_{1\flat})^{\sharp}(g)=|F|^{*}(g)-\theta_{2}(\bar{F}_{1})^{\sharp}(\rho^{*}(\mathsf{e}^{*}\mathrm{d}g))+(\theta_{2}-\theta_{1})(\bar{F}_{1})^{\sharp}(\mathsf{e}^{*}\mathsf{i}^{*}\mathrm{d}g)\\ -\theta_{2}\theta_{1}\frac{1}{2}[(\bar{F}_{1})^{\sharp}(\mathsf{e}^{*}\mathsf{i}^{*}\mathrm{d}g),(\bar{F}_{1})^{\sharp}(\mathsf{e}^{*}\mathsf{i}^{*}\mathrm{d}g)]~{},\end{gathered}$ (6.31)
and the lift of $F_{1\flat}^{\prime}$ to $F_{1}^{\prime}$ is simply
Let us again choose local coordinates $(m_{1}^{a},\bar{m}_{1}^{a},\xi^{\alpha},\bar{\xi}^{\alpha})$ for $\mathrm{T}[1]\mathfrak{g}[1]$ in a neighborhood $U\subset M$. The embedding $\mathsf{e}:M\rightarrow\mathsf{G}$ then allows us to identify $\mathsf{e}(U)\subset\mathsf{G}$ with $U\times V$, on which we use local coordinates $(m_{2}^{a},\bar{m}_{2}^{a},y^{\alpha},\bar{y}^{\alpha})$. In the fibered product $\mathrm{T}[1]\mathfrak{g}[1]\mathbin{{}_{\psi}\times_{\mathrm{d}\mathsf{t}}}\mathrm{T}[1]\mathsf{G}$ we need to identify $\bar{m}_{1}^{a}+\mathtt{r}^{a}_{\alpha}\xi^{\alpha}$ with $\bar{m}_{2}^{a}$ and $m_{1}$ with $m_{2}$. We will show later that
where $O_{2}$ represents functions of order at least 2 in its argument.
$\begin{split}(\bar{F}_{0})^{\sharp}(\xi^{\alpha})+\theta\mathrm{d}_{\mathcal{N}}(\bar{F}^{1})^{\sharp}(\xi^{\alpha})+\theta|F|^{*}(\omega^{\alpha}_{a\beta})F_{0}^{\sharp}(\bar{m}_{1}^{a})(\bar{F}_{1})^{\sharp}(\xi^{\alpha})+\theta|F|^{*}(\mathtt{f}_{\beta\gamma}^{\alpha})(\bar{F}_{1})^{\sharp}(\xi^{\beta})(\bar{F}_{0})^{\sharp}(\xi^{\gamma})~{},\end{split}$ (6.35)
It remains to prove (6.34). Recall that the Lie algebroid connection corresponding to the Lie groupoid connection is given by [CSS12, Thm. 3.11] (see also Proposition 2.12)
Firstly, we show that
Using the Baker–Campbell–Hausdorff formula, one can see that
Note that the left hand side is nothing but $\mathrm{d}R_{\mathsf{exp}(-\epsilon\gamma)}\sigma_{\mathcal{H}}\bigl{(}\mathsf{exp}(\epsilon\gamma_{x}),V_{x}\bigr{)}$ with $\epsilon\gamma=y^{\alpha}\bar{y}_{\alpha}$ and $V=\bar{m}$, which makes it clear that $\bar{U}(m,y)_{a\beta}^{\alpha}=-\omega_{a\beta}^{\alpha}$.
Recall that in Proposition 2.10 we showed that $\mathrm{Ad}(g^{-1},\gamma)=(g^{-1},0)\circ(\mathsf{e}_{\mathsf{t}(g)},\gamma)\circ\sigma_{\mathcal{H}}(g,\rho(\gamma))$. Using the Baker–Campbell–Hausdorff formula for Lie groupoids, one sees the dual of the map $\mathsf{p}_{\mathfrak{g}[1]}\circ\mathrm{Ad}\circ\mathsf{inv}^{\star}:\mathsf{t}^{*}\mathfrak{g}[1]\to\mathfrak{g}[1]$ takes the form
$=\left(m,0,-y+O_{2}(y),0\right)\circ\bigl{(}(m,\bar{m},y,\bar{y})-\sigma_{\mathcal{H}}(m,\bar{m},y,0)\bigr{)}~{}.$ (6.43)
With this, it is easy to see that the above map dualizes to